Problem 2.6 Annual deposits earning monthly interest Given: You deposit $10,000 each year at a monthly interest of 8%. Find: The amount you have in 10 years.
Answer #1

Amount= A= d[(1+i)^n - 1]/i

Where i= 8%/12= 0.08/12= 0.007

n= 12*10= 120

A= 10000[(1+0.007)^120 - 1]/0.007

= 1831291.6435